OpenStudy (anonymous):

Given the formula y=ax^(n-3), find n when y=192, a= 6, and x=2.

OpenStudy (australopithecus):

plug the numbers into the equation and solve for n you will need to use logarithms to solve for n using the rule log(x^(y)) = (y)log(x)

OpenStudy (anonymous):

i understand that it will be 192=(6)(2)^(n-3) but im not sure how solve it.

OpenStudy (australopithecus):

plug the numbers in and take the log of both sides

OpenStudy (anonymous):

I dont understand what log is.

OpenStudy (anonymous):

\[ 192 = 6 2^{n-3}\\ 32 = 2^{n-3} \] What power of 2 gives you 32? Read on \] 2^2= 4 2^3= 8 2^4=16 \] So \[ 2^{x} = 32 \] How much is x?

OpenStudy (anonymous):

To finish it n-3 = x n = 3+x